please help me solve this equation: 1/a+1/b=1/c. solve for b
Answer by ptaylor(2198)

1/a + 1/b=1/c multiply each term by abc
bc+ac=ab subtract bc from each side
bc-bc+ac=ab-bc collect like terms
ab-bc=ac or
b(a-c)=ac divide each side by (a-c)
b=ac/(a-c)--------------------ans

Does this help?---ptaylor